Train A travels 60mph
Train B travels 70mph
Train A reaches a spot at 6:15, Train B reaches same spot at 6:27, when will train B catch up to train A?

Answer by gonzo(654)

the trains will meet after they have both traveled the same distance.
train A has to do that distance in T minutes
train B has to do that distance in (T-12) minutes because train B started 12 minutes later (6:27 - 6:15).
-----
the equation for train A is T * 60 = D
the equation for train B is (T - 12) * 70 = D
since they both = D, then they are equal to each other.
T * 60 = (T - 12) * 70
solve this equation as follows:
T * 60 = (70 * T) - (70 * 12)
60*T = 70*T - 840
840 = 10*T
T = 84 minutes.
-----
in 84 minutes, train A will have traveled 84/60 * 60 = 84 miles.
in 84 minutes, train B will have traveled 72/60 * 70 = 84 miles.
-----
division by 60 was necessary to convert minutes to hours.